# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@
Abstract We propose a hybrid estimation procedure to estimate global fixed parameters and subject-specific random effects in a mixed fractional Black-Scholes model based on discrete-time observations. Specifically, we consider $N$ independent stochastic processes, each driven by a linear combination of standard Brownian motion and an independent fractional Brownian motion, and governed by a drift term that depends on an unobserved random effect with unknown distribution. Based on $n$ discrete time statistics of process increments, we construct parametric estimators for the Brownian motion volatility, the scaling parameter for the fractional Brownian motion, and the Hurst parameter using a generalized method of moments. We establish their strong consistency under the two-step regime where the observation frequency $n$ and then the sample size $N$ tend to infinity, and prove their joint asymptotic normality when $H\in\big(\frac{1}{2},\frac{3}{4}\big)$. Then, from one trajectory, we consistently estimate the random effects, using a plug-in approach, and we study their asymptotic behavior under different growth regimes of frequency and sample size. Finally, we construct a nonparametric estimator for the distribution function of these random effects using a Lagrange interpolation at Chebyshev-Gauss nodes based method, and we analyze its asymptotic properties as both $n$ and $N$ increase. We illustrate the theoretical results through a numerical simulation framework. We further demonstrate the efficiency performance of the proposed estimators in an empirical application to crypto returns data, analyzing five major cryptocurrencies to uncover their distinct volatility structures and heterogeneous trend behaviors.
@@ 1 Introduction
In many practical contexts, different forms of randomness influence the dynamics of a system. Short-term fluctuations or noise with independent increments are commonly modeled using the Wiener process. However, this process does not capture the long-range dependence observed in a wide range of real-world systems. To address this limitation, fractional Brownian motion (fBm) has been proposed, as a generalization of standard Brownian motion that exhibits long-range dependence, making it well-suited for modeling dynamic systems characterized by memory effects and self-similarity. The statistical inference for stochastic differential equations (SDE) models driven by a standard Brownian motion has been extensively studied in many papers and summarized in several books, see e.g, Bishwal (2008), Kutoyants (2004), Rao (1999) and references cited therein. As for SDE models governed by an fBm, there has been considerable interest in studying problems of statistical inference for these models and how to estimate the unknown parameters on which depends the model, see e.g. Mishura (2008), Rao (2011) and references therein. Although models governed by Wiener process or fBm capture either short-term noise or long-range dependence, respectively, many real-world systems exhibit both characteristics simultaneously. For instance, financial markets where asset prices exhibit both short-term volatility and long-term correlations. Similarly, models in telecommunications that have both rapid fluctuations and persistent network characteristics. Physiological signals often exhibit both fast, irregular fluctuations and underlying trends that persist over time.
@@
Parameter estimation problem for this process in a classical framework without random effects, has attracted several researchers and works, including Kukush et al. (2022) where two types of estimators were proposed, one based on power variations and the other based on the ergodic theorem and the generalized method of moments which have allowed a simultaneous estimation for all the parameters. A similar simultaneous estimation was investigated in Dufitinema et al. (2022), where the maximum likelihood approach is combined with numerical optimization methods. A particular case of model (1.1) where $\gamma=1$ and $H$, $\sigma$ are assumed to be known, was considered in Cai et al. (2016) where the authors constructed estimators for the drift parameter $\theta$. In Mishura et al. (2017), drift parameter estimation was investigated for a more general model driven by a Gaussian process with stationary increments. More recently, El Omari (2024) considered a more general version of model (1.1) without random effects, which he called $n$-th order mixed fractional Brownian motion with polynomial drift and constructed parametric estimators based on discrete observations for diffusion parameters and Hurst index, and on continuous observations for drift parameter and studied asymptotic properties of all estimators.
@@ 2.2 Approximation of a distribution function using Lagrange polynomials
Let $Z_{1},\ldots,Z_{N}$ be a sequence of independent and identically distributed (in short i.i.d) random variables with a common unknown distribution function $G$ supported on [1,1][-1,1]. The ordinary nonparametric estimator of order $m\geq 1$ for the distribution function $G$ obtained by using an approximation based on Lagrange polynomials with Chebyshev Gauss points, is defined for all $x\in[-1,1]$ as follows
@@ 3. $(A3)$
  • Assumption $(A1)$ is standard for applying the multivariate central limit theorem (CLT) and the delta method rigorously. It ensures that the fourth moment of the random effects is finite, which is necessary to control the asymptotic variance of the second-order sample moments $\bar{V}_{n,N}$ and, consequently, of the estimators $\widehat{\Theta}$.
  • Assumptions $(A2)$ and $(A3)$ are classical in nonparametric estimation theory using interpolation methods. The $C^{2}$ smoothness of $F$ allows for a second-order Taylor expansion, which is crucial for evaluating the bias of the Lagrange interpolant estimator $\hat{F}_{m,n,N}$ in Proposition 5. The boundedness of $f$ and $f^{\prime}$ is needed to control the behavior of the distribution function and the estimation error uniformly across its domain, including the boundaries.
@@ 3.1 Estimation of model parameters and random effects
The construction of all parameter estimators and the study of their asymptotic behavior are based on the following ergodic result.
@@
In the next proposition, we construct and establish the asymptotic properties of $\theta_{i}$ for each fixed $i=1,\ldots,N$.
@@
Based on these increments, for each $i=1,\ldots,N$, we introduce the following per-subject statistics that will be used in the construction of the parameter estimators.
@@
The estimators of the global parameters and their asymptotic properties will be derived from the asymptotic behavior of the above empirical statistics. We first establish their almost sure convergence and asymptotic normality based on the ergodic property proved in Lemma 1. In the next proposition, we establish the almost sure convergence of the averaged statistics $\bar{U}_{n,N}:=\Big(\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}\Big)^{\top}.$
@@
In the next proposition, we establish the asymptotic normality of the empirical statistics.
@@
where $\Sigma:=\mathrm{Var}\big(\theta_{i}^{2}\big)\,\mathcal{H}\,\mathcal{H}^{\top}$ with $\mathcal{H}=\big(1,h^{2},h^{2},4h^{2}\big)^{\top}$.
@@
Under Assumption $(A1)$, $\widehat{\Theta}:=\left(\hat{H},\hat{\gamma}^{2},\hat{\sigma}^{2}\right)^{\top}$ is a strongly consistent estimator of $\Theta$, that is
@@
The next theorem concerns the joint asymptotic normality of $\widehat{\Theta}_{N}$.
@@
$\sqrt{N}\,\big(\widehat{\Theta}_{N}-\Theta\big)\ \xrightarrow{d}\ \mathcal{N}\!\left(0,\;J\,\Sigma\,J^{\!\top}\right),$
where $J$ is the Jacobian
$J=\left.\frac{\partial\widehat{\Theta}}{\partial\bar{U}_{n,N}}\right|_{U_{\infty}}=\begin{pmatrix}\left.\dfrac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\\[5.69054pt] \left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\\[5.69054pt] \left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\end{pmatrix}\!$,
where the derivatives of $\hat{H}$ are given as follows
@@
the derivatives of $\hat{\gamma}^{2}$ are expressed as
@@
and finally the derivatives of $\hat{\sigma}^{2}$ are given by
@@
The strong consistency and asymptotic normality of $\widehat{\phi}_{i}$, $i=1,\ldots,N$, are established in the following proposition.
@@
    As $n,N\to\infty$ sequentially, the estimator $\widehat{\phi}_{i}$ is strongly consistent, that is, for $i=1,\ldots,N$, $\widehat{\phi}_{i}\ \xrightarrow{a.s}\ \phi_{i}.$
@@
    1. (i)
      If $n,N\to\infty$ sequentially, we obtain
@@
      where $v_{\sigma}^{2}$ is the asymptotic variance of $\hat{\sigma^{2}}$.
    2. (ii)
      If $n,N\to\infty$ simultaneously, we obtain three critical asymptotic cases:
      • If $\frac{n^{1-H}}{\sqrt{N}}\to c\in(0,\infty)$,  $n^{1-H}\Big(\widehat{\phi}_{i}-\phi_{i}\Big)\xrightarrow{d}\mathcal{N}\Big(0,v_{\theta}^{2}+\tfrac{c^{2}}{4}v_{\sigma}^{2}\Big)$.
      • If $\frac{n^{1-H}}{\sqrt{N}}\to 0$,  $n^{1-H}\Big(\widehat{\phi}_{i}-\phi_{i}\Big)\xrightarrow{d}\mathcal{N}\Big(0,v_{\theta}^{2}\Big)$,  with $v_{\theta}^{2}=\gamma^{2}h^{2H-2}.$
      • If $\frac{n^{1-H}}{\sqrt{N}}\to\infty$,  $\sqrt{N}\Big(\widehat{\phi}_{i}-\phi_{i}\Big)\xrightarrow{d}\mathcal{N}\Big(0,\tfrac{1}{4}v_{\sigma}^{2}\Big)$.
Remark 1.
To study the asymptotic behavior of the global parameter estimator $\widehat{\Theta}$, we considered the two-step regime where $n\to\infty$ first and then $N\to\infty$, while for the random effects $\phi_{i}$ we investigated both the two-step and simultaneous regimes. The reason is technical. Indeed, the global parameter estimators are based on the averaged statistics $(\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N})$, whose limits require first letting $n\to\infty$ to use ergodicity within each subject’s trajectory and then $N\to\infty$ to apply the SLLN across subjects. If instead $n$ and $N$ increase simultaneously, the SLLN for $(\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N})$ depends on their relative growth rates: if $N$ grows too fast, the per-subject statistics may not be close enough to their expectations, which biases the averages. For this reason, the two-step regime is the natural framework for the global parameters. By contrast, the estimators of the random effects $\phi_{i}$ are based on per-subject statistics and involve $\hat{\sigma}^{2}$. In this case, the asymptotic error can be decomposed into an $n$-dependent part and an $N$-subject averaging part, which justifies studying both regimes.
In the next subsection, we proceed to estimate the distribution $F$ of random effects based on the estimated sample $\widehat{\phi}_{1},\ldots,\widehat{\phi}_{N}$.

### v4

@@
Abstract We propose a hybrid estimation procedure to estimate global fixed parameters and subject-specific random effects in a mixed fractional Black-Scholes model based on discrete-time observations. Specifically, we consider $N$ independent stochastic processes, each driven by a linear combination of standard Brownian motion and an independent fractional Brownian motion, and governed by a drift term that depends on an unobserved random effect with unknown distribution. Based on $n$ discrete time statistics of process increments, we construct parametric estimators for the Brownian motion volatility, the scaling parameter for the fractional Brownian motion, and the Hurst parameter using a generalized method of moments. We establish their strong consistency under the two-step regime where the observation frequency $n$ and then the sample size $N$ tend to infinity, and prove their joint asymptotic normality when $H\in\big(\frac{1}{2},\frac{3}{4}\big)$. Then, using a plug-in approach, we consistently estimate the random effects, and we study their asymptotic behavior under the same sequential asymptotic regime. Finally, we construct a nonparametric estimator for the distribution function of these random effects using a Lagrange interpolation at Chebyshev-Gauss nodes based method, and we analyze its asymptotic properties as both $n$ and $N$ increase. We illustrate the theoretical results through a numerical simulation framework. We further demonstrate the efficiency performance of the proposed estimators in an empirical application to crypto returns data, analyzing five major cryptocurrencies to uncover their distinct volatility structures and heterogeneous trend behaviors.
@@ 1 Introduction
In many practical contexts, different forms of randomness influence the dynamics of a system. Short-term fluctuations or noise with independent increments are commonly modeled using the Wiener process. However, this process does not capture the long-range dependence observed in a wide range of real-world systems. To address this limitation, fractional Brownian motion (fBm) has been proposed, as a generalization of standard Brownian motion that exhibits long-range dependence, making it well-suited for modeling dynamic systems characterized by memory effects and self-similarity. The statistical inference for stochastic differential equations (SDE) models driven by a standard Brownian motion has been extensively studied in many papers and summarized in several books, see e.g, Bishwal (2008), Kutoyants (2004), Rao (1999) and references cited therein. As for SDE models governed by an fBm, there has been considerable interest in studying problems of statistical inference for these models and how to estimate the unknown parameters on which the model depends, see e.g. Mishura (2008), Rao (2011) and references therein. Although models governed by the Wiener process or fBm capture either short-term noise or long-range dependence, respectively, many real-world systems exhibit both characteristics simultaneously. For instance, financial markets where asset prices exhibit both short-term volatility and long-term correlations. Similarly, models in telecommunications that have both rapid fluctuations and persistent network characteristics. Physiological signals often exhibit both fast, irregular fluctuations and underlying trends that persist over time.
@@
Parameter estimation problem for this process in a classical framework without random effects, has attracted several researchers and works, including Kukush et al. (2022) where two types of estimators were proposed, one based on power variations and the other based on the ergodic theorem and the generalized method of moments which have allowed a simultaneous estimation for all the parameters. A similar simultaneous estimation was investigated in Dufitinema et al. (2022), where the maximum likelihood approach is combined with numerical optimization methods. A particular case of model (1.1) where $\gamma=1$ and $H$, $\sigma$ are assumed to be known, was considered in Cai et al. (2016) where the authors constructed estimators for the drift parameter $\theta$. In Mishura et al. (2017), drift parameter estimation was investigated for a more general model driven by a Gaussian process with stationary increments. More recently, El Omari (2024) considered a more general version of model (1.1) without random effects, which he called $n$-th order mixed fractional Brownian motion with polynomial drift and constructed parametric estimators based on discrete observations for diffusion parameters and Hurst index, and on continuous observations for drift parameter, and studied asymptotic properties of all estimators.
@@ 2.2 Approximation of a distribution function using Lagrange polynomials
Let $Z_{1},\ldots,Z_{N}$ be a sequence of i.i.d random variables with a common unknown distribution function $G$ supported on [1,1][-1,1]. The ordinary nonparametric estimator of order $m\geq 1$ for the distribution function $G$ obtained by using an approximation based on Lagrange polynomials with Chebyshev Gauss points, is defined for all $x\in[-1,1]$ as follows
@@ 3. $(A3)$
  • Assumption (A1) is needed for studying the asymptotic properties of our model parameters via the delta method and the multivariate central limit theorem. It enables the control of squared terms that appear in second-order statistics, which we will use in our parametric estimation procedure.
  • Assumptions $(A2)$ and $(A3)$ are classical in nonparametric estimation theory using interpolation methods. The $C^{2}$ smoothness of $F$ allows for a second-order Taylor expansion, which is crucial for evaluating the bias of the Lagrange interpolant estimator $\widehat{F}_{m,n,N}$ in Proposition 6. The boundedness of $f$ and $f^{\prime}$ is needed to control the behavior of the distribution function and the estimation error uniformly across its domain, including the boundaries.
@@ 3.1 Estimation of model parameters and random effects
The construction of all parameter estimators and the study of their asymptotic behavior are based on the following ergodic result, which will be proved in the proofs section.
@@
Based on this lemma, we present and study a natural estimator for the parameter $\hat{\theta}_{i}$ in the following proposition.
@@
Now, based on the estimator $\hat{\theta}_{i}$ and the increments of $Y^{i}$, we introduce the following per-subject statistics that will be used in the construction of the parameter estimators.
@@
The estimators of the global parameters and their asymptotic properties will be derived from the asymptotic behavior of the above empirical statistics. In the next proposition, we establish the almost sure convergence of the statistics $\bar{U}_{n,N}:=\Big(\bar{V}_{n,N},\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N}\Big)^{\top}.$
@@
Remark 1.
The sequential asymptotic regime $n\to\infty$ first, then $N\to\infty$ is essential for the consistency and asymptotic normality of the subsequent global parameter estimators $\widehat{\Theta}=(\widehat{H},\widehat{\gamma}^{2},\widehat{\sigma}^{2})^{\top}$. The reason is technical. Indeed, these estimators are constructed from the averaged statistics $(\bar{\xi}_{n,N},\bar{\eta}_{n,N},\bar{\zeta}_{n,N})$, whose limits require first letting $n\to\infty$ to use ergodicity within each subject’s trajectory and then $N\to\infty$ to apply the SLLN across subjects. If instead $n$ and $N$ increase simultaneously at comparable rates, the individual statistics $\xi^{i}_{n}$, $\eta^{i}_{n}$, and $\zeta^{i}_{n}$ may not be sufficiently close to their limits $\xi^{i}_{\infty}$, which introduces bias in the averaged statistics. Practically, the sequential regime corresponds to the realistic scenario where we collect dense observations ($n$ large) for each trajectory to accurately estimate its specific characteristics, and then aggregate information across many trajectories ($N$ large) to estimate population-level parameters. For this reason, all asymptotic results are investigated under the sequential regime.
The next proposition establishes the asymptotic normality of the empirical statistics.
@@
$\Sigma:=\mathrm{Var}\big(\theta_{i}^{2}\big)\,\mathcal{H}\,\mathcal{H}^{\top}\quad\text{with}\quad\mathcal{H}=\big(1,h^{2},h^{2},4h^{2}\big)^{\top}.$ (3.8)
@@
Under Assumption $(A1)$, $\widehat{\Theta}:=\left(\widehat{H},\widehat{\gamma}^{2},\widehat{\sigma}^{2}\right)^{\top}$ is a strongly consistent estimator of $\Theta$, that is
@@
The next theorem concerns the asymptotic normality of $\widehat{\Theta}_{N}$.
@@
$\sqrt{N}\,\big(\widehat{\Theta}_{N}-\Theta\big)\ \xrightarrow{d}\ \mathcal{N}\!\left(0,\;\mathcal{C}\right),$
where the asymptotic covariance matrix $\mathcal{C}$ is given by $\mathcal{C}:=J\,\Sigma\,J^{\!\top}$ with $\Sigma$ is the matrix (3.8) and the Jacobian matrix $J$ is defined as follows
$J:=\left.\frac{\partial\widehat{\Theta}}{\partial\bar{U}_{n,N}}\right|_{U_{\infty}}=\begin{pmatrix}\left.\dfrac{\partial\widehat{H}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{H}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\\[5.69054pt] \left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\gamma}^{2}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\\[5.69054pt] \left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{V}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\xi}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\eta}_{n,N}}\right|_{U_{\infty}}&\left.\dfrac{\partial\widehat{\sigma}^{2}}{\partial\bar{\zeta}_{n,N}}\right|_{U_{\infty}}\end{pmatrix}\!$,
where the derivatives of $\widehat{H}$ are given as follows
@@
the derivatives of $\widehat{\gamma}^{2}$ are expressed as
@@
and finally the derivatives of $\widehat{\sigma}^{2}$ are given by
@@
Remark 2.
The restriction $H\in\left(\frac{1}{2},\frac{3}{4}\right)$ ensures that the statistics vector $U_{n,N}$ satisfies a Gaussian central limit theorem under the scaling $\sqrt{N}$. This stems from the Breuer-Major theorem (Theorem A.2), which requires the condition
$\sum_{k\in\mathbb{Z}}|\rho(k)|^{2}<\infty,$ (3.12)
where $\rho(k)$ is the autocovariance function defined in (7.1). Using the asymptotic behavior of $\rho_{H}(k)$, which is given by $\rho_{H}(k)\sim h^{2H}H(2H-1)k^{2H-2}$ as $k\to\infty$, we obtain
$\sum_{k=1}^{\infty}|\rho_{H}(k)|^{2}\sim Ch^{4H}\sum_{k=1}^{\infty}k^{4H-4},$ (3.13)
which converges if and only if 4H4<14H-4<-1, i.e., $H<\frac{3}{4}$. At the critical value $H=\frac{3}{4}$, the asymptotic behavior of functionals of fractional Gaussian noise typically changes: for $H<\frac{3}{4}$, one obtains a Gaussian limit under standard normalizations, whereas for $H\geq\frac{3}{4}$ the limit becomes non-Gaussian and typically involves Rosenblatt-type distributions (see for instance Dobrushin and Major (1979), Kubilius et al. (2017), Rosenblatt (1952), Taqqu (1975)). The case $H\geq\frac{3}{4}$ requires different techniques involving multiple Wiener-Itô integrals and Malliavin calculus, which is beyond the scope of this paper.
@@
Remark 3.
The estimators $\widehat{\phi}_{1},\ldots,\widehat{\phi}_{N}$ are not independent because they all use the same plug-in estimator $\widehat{\sigma}^{2}$. Under our sequential regime, we show that $\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|\overset{a.s}{\longrightarrow}0$ (see equation (7.32)).
Hence, the dependence does not affect the uniform consistency. For distributional limits, the common error $\frac{1}{2}(\widehat{\sigma}^{2}-\sigma^{2})=O_{\mathbb{P}}(N^{-\frac{1}{2}})$ must be controlled at the corresponding rate; this is done explicitly in the proofs of Theorems 3-4 and Proposition 7).
The following Proposition provides the strong consistency and asymptotic normality of $\widehat{\phi}_{i},i=1,\dots,N$, for the sequential regime ($n\to\infty$ then $N\to\infty$).
@@
    As $n,N\to\infty$ sequentially, the estimator $\widehat{\phi}_{i}$ is strongly consistent, that is, for each fixed $i=1,\ldots,N$, $\widehat{\phi}_{i}\ \xrightarrow{a.s}\ \phi_{i}.$
@@
    As $n,N\to\infty$ sequentially, we obtain
@@
    where $v_{\sigma}^{2}$ is the asymptotic variance of $\widehat{\sigma^{2}}$.
In the next subsection, we proceed to estimate the distribution $F$ of random effects based on the estimated random effects $\widehat{\phi}_{1},\ldots,\widehat{\phi}_{N}$.